Consider the series:


"\u221e"

"\u03a3" [ "(x^2 + 9)" / 25 "]^a"

"a = 1"


The values of "y" of which this series converges form an internal "(-k, k)" .

Give the value for "k"


1
Expert's answer
2021-06-02T16:35:48-0400

It is given that,

"\displaystyle\sum_{a=1}^\infin\lbrack\frac{(x^2+9)}{25}\rbrack^a=\displaystyle\sum_{a=1}^\infin\lbrack M_a\rbrack"


"\implies M_a=\lbrack\frac{(x^2+9)}{25}\rbrack^a"


By the ratio test,

"L=\lim_{a\rightarrow \infty} \vert\frac{M_{a+1}}{M_a}\vert"

"L=\lim_{a\rightarrow \infty }\vert\frac{\lbrack\frac{ (x^2+9)}{25}\rbrack ^{a+1}}{\lbrack \frac{(x^2+9)}{25}\rbrack^a}\vert"


"L=lim_{ a\rightarrow \infty}\space \vert\lbrack\frac{(x^2+9)}{25}\rbrack\vert"


"L=\vert \lbrack\frac{(x^2+9)}{25}\rbrack\vert"


For interval of convergence,

L<1

"\vert\frac{(x^2+9)}{25}\vert<1"


"\frac{x^2+9}{25}<1"


"x^2+9<25\\x^2+9-25<0\\x^2-16<0\\(x-4)(x+4)<0\\-4<x<4"


Hence the interval of convergence is "x\in(-4,4)"


and "k=4"